Demand for oil changes at​ Garcia's Garage has been as​ follows:

                                                                 

Month

Number of Oil Changes

January

43

February

51

March

63

April

50

May

60

June

42

July

68

August

72

a. Use simple linear regression analysis to develop a forecasting model for monthly demand. In this​ application, the dependent​ variable, Y, is monthly demand and the independent​ variable, X, is the month. For​ January, let

Xequals=​1;

for​ February, let

Xequals=​2;

and so on.

The forecasting model is given by the equation

Yequals=+X.

​(Enter your responses rounded to two decimal​ places.)
Answer #1


PERIOD

DEMAND

X

Y

XY

X^2

1

43

1

43

43

1

2

51

2

51

102

4

3

63

3

63

189

9

4

50

4

50

200

16

5

60

5

60

300

25

6

42

6

42

252

36

7

68

7

68

476

49

8

72

8

72

576

64

SIGMA

36

449

2138

204


INTERCEPT(B0) = (SIGMA(Y) * SIGMA(X^2) - SIGMA(X) * SIGMA(XY)) / (N * SIGMA(X^2) - SIGMA(X)^2)

INTERCEPT = (449 * 204) - (36 * 2138) / ((8 * 204) - 36^2) = 43.54

SLOPE(B1) = ((N * SIGMA(XY)) - (SIGMA(X) * SIGMA(Y))) - (N * SIGMA(X^2) - SIGMA(X)^2)

SLOPE = ((8 * 2138) - (36 * 449) / ((8 * 204) - 36^2) = 2.8

EQUATION

Y = INTERCEPT + SLOPE(X)

Y = 43.54 + (2.8 * (X))

FOR X = 9(SEPTEMBER), FORECAST = 43.54 + (2.8 * 9) = 68.74